uestion 66 In March, Kelly Company had the following unit production costs: materials $12 and conversion costs $7. On March 1, it had no work in process. During March, Kelly transferred out 22,500 units. As of March 31, 3,800 units that were 48% complete as to conversion costs and 100% complete as to materials were in ending work in process. Compute the total units to be accounted for. Total units LINK TO TEXT Compute the equivalent units of production. Materials Conversion costs The equivalent units of production LINK TO TEXT Prepare a cost reconciliation schedule, including the costs of materials transferred out and the costs of materials in process. Cost Reconciliation Schedule Costs accounted for Transferred out $ Work in process, March 31 Materials $ Conversion costs Total costs $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

Answer 1:

Total units to be accounted for = 26,300

Workings:

Total units to be accounted for = Units transferred out + Units in Ending work in process = 22,500 + 3,800 = 26,300

Answer 2:

Answer 3:

Given:

Conversion costs per unit of production = $7

Materials costs per unit of production = $12

Costs accounted for:

Transferred out = 22500 * (12 + 7) = $427,500

Work in process, March 31:

Materials = 3800 * 12 = $45,600

Conversion costs =1824 * 7 = $12,768

Total work in process costs = $58,368

Total costs accounted for = 427500 + 58368 = $485,868